(2x+5)2=9
q2-2(7q+16)=0

Answer by checkley75(3666)

(2x+5)2=9
4x+10=9
4x=9-10
4x=-1
x=-1/4 answer
proof
(2*-1/4+5)2=9
(-1/2+5)2=9
4.5*2=9
9=9
----------------------------------------------------
q^2-2(7q-16)=0
q^2-14q+32=0
(q-16)(q+2)=0
q-16=0
q=16 answer.
q=2=0
Q=-2 answer.